OpenStudy (anonymous):

Explain how you calculate the restricted value(s) of x for the equation the fraction the binomial x plus 3 over the binomial x plus 5 plus the fraction x over 6 equals the fraction 5 over 6 and what those value(s) represent. equation included.

OpenStudy (anonymous):

\[\frac{ x+3 }{ x+5 }+\frac{ x }{ 6 }=\frac{ 5 }{ 6 }\]

OpenStudy (anonymous):

@Hero ?

hero (hero):

The restricted value of x occurs where the denominator cannot equal zero, so to find this restricted value solve x + 5 = 0.

OpenStudy (anonymous):

x=-5

hero (hero):

yes, correct.

OpenStudy (anonymous):

then it can be anything except -5?

hero (hero):

Yes, correct.

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

one more question. would you use cross-multiplication or LCD to solve it?

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

@hero?

hero (hero):

If you know what you're doing, you could use either method.

OpenStudy (anonymous):

which one is easier?

hero (hero):

That depends on how familiar you are with either method. But in general, the easier method tends to be the method that requires the least amount of steps to solve.

OpenStudy (anonymous):

which is...?

hero (hero):

I'd say it's different for different people. Me personally, I would use the cross multiplication method after putting the equation in proper form.

OpenStudy (anonymous):

how would you do that again?

hero (hero):

There's a particular reason for this. If you subtracted 5/6 from both sides you would get: \[\frac{x + 3}{x + 5} + \frac{x}{6} - \frac{5}{6} = 0\] Which becomes: \[\frac{x + 3}{x + 5} + \frac{x - 5}{6} = 0\] Then after subtracting (x-5)/6 from both sides: \[\frac{x + 3}{x + 5} =- \frac{x - 5}{6}\] You would cross multiply to get \[6(x + 3) = -(x - 5)(x + 5)\] The right hand side represents a difference of squares: \[6x + 18 = -(x^2 - 25)\] \[6x + 18 = -x^2 + 25\] From there, you would re-arrange it into a quadratic equation then solve.

OpenStudy (anonymous):

how would it look once it was finished?

OpenStudy (anonymous):

@Hero

hero (hero):

Have you attempted to finish solving it?

OpenStudy (anonymous):

yesh but you cant put it into quadratic equations because it doesnt have a ^2

hero (hero):

It doesn't? Are you sure about that?

OpenStudy (anonymous):

oh sorry. didnt see that! derp

OpenStudy (anonymous):

-x^2-6x+7 right?

hero (hero):

Don't forget to set it equal to zero. You want to make sure the leading term is non-negative. It is easier to factor this way. What you do is this: \[-x^2 - 6x + 7 = 0\] Factor out -1: \[-(x^2 + 6x - 7) = 0\] Then divide both sides by -1 \[x^2 + 6x - 7 = 0\] Now you should be able to factor relatively easy

OpenStudy (anonymous):

ok thanks! gtg

hero (hero):

Okay, good luck with everything :)